# DOSA-MO: Dual-stage Optimizer for Systematic overestimation Adjustment in Multi-Objective problems improves biomarker discovery

### 27 Dec 2023 v1

@@ Abstract
The challenge in biomarker discovery and validation using machine learning from omics data lies in the abundance of molecular features but scarcity of samples. Most machine learning-based feature selection methods necessitate of hyperparameter tuning, typically performed by evaluating numerous alternatives on a validation set. Every evaluation has a performance estimation error and when the selection takes place between many models the best ones are almost certainly overestimated. Biomarker identification is a typical multi-objective problem with trade-offs between the predictive ability and the parsimony in the number of molecular features. Genetic algorithms are a popular tool for multi-objective optimization but they evolve numerous solutions and are prone to overestimation. Methods have been proposed to reduce the overestimation after a model has already been selected in single-objective problems, but to the best of our knowledge no algorithm existed that was capable of reducing the overestimation during the optimization, leading to a better model selection, or that had been applied in the more general domain of multi-objective problems. We propose DOSA-MO, a novel multi-objective optimization wrapper algorithm that learns how the original estimation, its variance, and the feature set size of the solutions predict the overestimation, and adjusts the expectation of the performance during the optimization, improving the composition of the solution set. We verify that DOSA-MO improves the performance of a state-of-the-art genetic algorithm on left-out or external sample sets, when predicting cancer subtypes and/or patient overall survival, using three transcriptomics datasets for kidney and breast cancer. Since to the best of our knowledge there was no measure of estimation error for multi-objective solution sets, we propose two such measures. According to both of them, DOSA-MO provides more realistic performance estimates in all the considered scenarios.
Data and source code availability. The gene expression data used in this study is from public repositories. The source code and detailed numerical results will be available in a public server after peer review.
@@ 1 Introduction
Molecular biomarker discovery with machine learning (ML) is usually limited by data that includes many features but few samples. This renders the models prone to overfitting and the evaluation prone to estimation error. Many ML approaches involve hyperparameter tuning and feature selection, with model selection based on the performance measured on a validation set within a training, validation, and test paradigm. Comparing many models and choosing the best is prone to overestimation, i.e. the difference between the performance measured on the validation set and the real performance, a.k.a. “winners course”. The models that fit the noise present in the validation set are advantaged, a phenomena sometimes referred as overfitting on the validation set [1]. This is also exacerbated by data scarcity. In the quest to reach higher accuracies, increasing the number of hyperparameter configurations may lead to selected models with better performance on the validation set, but a corresponding increase in overestimation may determine diminishing or even negative improvements on the test set.
In biomarker selection both the predictive performance (in terms of disease type, survivability, etc.) and the practical applicability are crucial factors. A biomarker that includes a small number of molecular features requires less resources when applied clinically, thus might be preferred over a slightly more accurate but more expensive or time consuming alternative.
@@
Characterizing all the best compromises (or trade-offs) between predictive value and feature set size is a multi-objective (MO) optimization problem [2], and it can be solved by means of MO feature selection (MOFS) techniques [3, 4, 5, 6]. These techniques aim to identify not just a single best solution, as in single-objective (SO) problems, but rather a Pareto front of solutions. This front is the set of optimal solutions that illustrate the trade-offs between different objectives in MO optimization. However, all candidate solutions (or biomarker models selected by the employed MO feature selection technique) are evaluated on the validation set, which can result in the overestimation of the performance of the selected models.
K-fold cross-validation (CV) is the most common methodology for ML assessment and its benefits and limitations have received much attention in the SO case [7, 8, 9]. K-fold CV returns a model trained on all the available samples and an estimation of its performance computed by averaging $k$ CV results. The obtained model performance is usually underestimated when only one hyperparameter configuration is used. However, it tends to be overestimated when multiple configurations are evaluated, and the model with the best performance is returned [1].
While some methods have been proposed to improve the estimation of the performance when k-fold CV is applied for model selection, their design and application has been limited to SO problems that are a special subset of the more general MO problem. To the best of our knowledge, no attempt has been made to improve the performance estimation in MO problems like the identification of the best biomarkers in high-dimensional omics data.
MO problems may require more model evaluations when many good trade-offs exist to be discovered. It may be argued that this offers the possibility to model differences in overestimation in function of the position in the non-dominated front.
Limitedly to SO problems, approaches exists to improve the performance estimation of a selected model, but the model is still selected using the unadjusted estimation, so there is no improvement in the model selection. In the model selection process, where solutions are ranked based on a model evaluation metric, simply subtracting a constant from this metric will not affect the ranking and, therefore, the model selection. If instead the adjustment takes into account some characteristic of the models, such as the variance in performance during the validation phase or the feature set size, it might change their order. It may be argued that such an adjustment could improve the model selection if the model selection would use the adjusted estimations.
While the quality of the performance estimation can be easily measured in SO problems as the absolute difference between the performance expected by the algorithm and the performance measured on new samples, to the best of our knowledge no such measure exists for MO problems, where the results are solution sets. Here we propose two new measures for the quality of the performance estimation.
In Cattelani et al., Fig. 1 and Supplementary Fig. 1 are examples of solution set performance, resulting after executing MO optimization for breast cancer subtypes biomarker discovery [5]. The optimizer explored the best trade-offs between balanced accuracy and number of features, and used 3-fold CV internally to evaluate the solutions. The overestimation (difference between the “inner cv” and “test” dots in the figures) of the balanced accuracy increases as the number of features increases. Analogously, the overestimation increases when the estimation increases. Correlations like these may suggest that characteristics of the solutions might be predictive of the overestimation.
To the best of our knowledge, no previous work experimented the effectiveness of methods for mitigating the overestimation in ML applications with 2 or more objectives. Additionally, no previous work applied the adjustment to the performance estimation prior to the selection of the solution set, thus affecting the selection. Consequently, this approach is also yet to be explored in biomarker discovery within large-scale omics data, where adjusting the expected performance of evaluated biomarker models across multi-cohort datasets could lead to a more robust and reliable biomarker selection.
We present a new MO optimization algorithm that wraps any other MO optimizer to produce temporary solutions. These solutions are then employed to train regression models, which are tasked with predicting performance overestimation. Subsequently, the algorithm executes a wrapped MO optimizer once more, but with a key distinction: it now utilizes the regression models to deliver adjusted fitness evaluations, enhancing the model selection process. The final resulting models are then trained on all the available samples, thus no samples are lost in order to compute their expected performance on new data.
In our benchmarking study, which concentrates on selecting cancer biomarkers from transcriptomic data, we have experimentally verified that incorporating information from the original performance estimation, its variance, and the size of the feature set can enhance both the performance and its estimation. This improvement is evident when predicting cancer subtypes and patient overall survival, particularly on test sets that are either excluded from the original dataset or entirely external.
Tsamardinos et al. [10] compare double CV, the Tibshirani and Tibshirani method [11], and nested CV in their ability to improve the estimation of the fitness for SO problems. These algorithms adjust the fitness estimation but do not affect the model selection, and the model that results is the same that is selected with simple k-fold CV with hyper-parameter optimization.
Automated ML (AutoML) tools try numerous combinations of models and hyperparameters and return the best performing model and an estimation of its performance. In Tsamardinos et al. six AutoML tools are compared [12]. Of these, only one has a predictive performance estimation strategy that adjusts for multiple model validations (limitedly to SO problems), while the majority of the tools have the necessity to withold a test set for an unbiased estimation of the performance of the winning model, thus loosing samples from the final model training. Differently from our proposed algorithm, none of the cosidered AutoML tools applies a performance adjustment in MO problems, and no considered tool uses the performance adjustment during the optimization for improving the selection of the winning model. To the best of our knowledge our approach is the first that can be applied to the more general MO problems and that improves the selection of the solutions in addition to their fitness estimation.
@@ 2 Methods
The DOSA-MO algorithm wraps a MO optimizer, serving two main purposes: improving the estimate of the performance of the solutions, and increase the overall performance of the solution set (set of biomarker models in our case study). The data-driven approach complicates the measurement of these achievements, as CV yields varied performance metrics. These include the performance anticipated by the optimizer, based on training data, and the actual performance measured on left-out samples post-optimization. To our knowledge, there is not an established metric for the error in evaluating the predictive performance of solutions in MO problems within a CV framework. In single-objective scenarios, one might simply assess the absolute difference between the fitness expected from training data and the fitness observed on new data. However, in MO scenarios, it is crucial to consider each solution’s contribution to the Pareto front. In this section, we introduce two novel metrics to measure this estimation error in MO CV setups.
Most importantly, we introduce a novel wrapper algorithm to improve the precision of the fitness estimation and the overall performance in MO optimization and prove its effectiveness in feature selection for biomarker discovery. Our wrapper algorithm follows a two-stage process using inner MO optimizers. In our case study, these optimizers are represented by genetic algorithms (GAs) specifically designed for MO problems and paired with supervised ML algorithms, such as the Gaussian Naïve Bayes (NB) classifier to distinguish different cancer subtypes or the Cox Proportional-Hazards Model (Cox) for survival analysis. In the first stage, regression-based models are trained to estimate the overestimation, drawing on characteristics of the solutions selected by an MO optimizer. Key characteristics include the original performance, its variance, assessed using only training data, and the size of the feature set. Subsequently, in the second stage, these trained regression models are employed to enhance the evaluation of solutions by an inner MO optimizer. This step aims to refine the selection of the solutions, thereby improving the overall results of the process.
We have benchmarked our algorithm, DOSA-MO, in the context of MOFS for cancer biomarker discovery. This was done using gene-based transcriptomic datasets from The Cancer Genome Atlas (TCGA) project [13]. Specifically, our goal is to identify biomarkers for categorizing cancer subtypes and survival prediction in kidney and breast cancer patients through a MOFS approach. For breast cancer, an additional cohort from The Sweden Cancerome Analysis Network - Breast (SCAN-B) [14] serves as an external validation set. Moreover, as inner MO optimizer, we use a novel modification of Non-dominated Sorting Genetic Algorithm III (NSGA3) [15], NSGA3 with clone-handling method and specialized mutation operator (NSGA3-CHS), with NB or Support Vector Machine (SVM) as inner classifiers, and Cox as inner survival model. We compare the unadjusted optimizer with adjustments by 5 different regression models, using both internal-external CV and external validation.
2.1 Measuring overestimation in multi-objective problems
We propose two methods to measure the error of the performance estimates for the solutions to MO problems: MO performance error and Pareto delta.
2.1.1 Multi-objective performance error
We define the MO performance error $E_{\upsilon}$ starting from the HyperVolume (HV) computed on the train performance ($H_{\iota}$) and the Cross HyperVolume (CHV) ($H_{\upsilon}$). $H_{\iota}$ and $H_{\upsilon}$ have been formally defined by Cattelani et al. ([6]). Since $H_{\upsilon}$ is a family of functions, with the specific instantiation that depends on the user provided function $\upsilon$, $E_{\upsilon}$ is a family of functions too.
Where $X$ encodes the performance of the solutions evaluated on the train data, and $X^{\prime}$ the performance evaluated on the test data. In our case study we use the same instantiation of $\upsilon$ as in Cattelani et al.: the function $\lambda$ ([6]).
2.1.2 Pareto delta
The MO performance error is a metric that derives straightforwardly from the train HV and the CHV. This metric does not pinpoint the specific origins of differences between these two measures. Even though the CHV inherently considers the differences between train and test performance inside its formulation, increases in performance in part of the solutions can even out with decreases in another part, leading to a reduced performance error measurement despite these differences between train and test performance that have opposite sign. To address this issue, we introduce an additional measure called Pareto delta. It has the property of being equal to 00 only when there is no difference in performance between train and test data for all elements of the solution set.
To calculate the Pareto delta, we first sum up the absolute error in fitness estimation for each solution, multiplying it by the derivative of the HV concerning that specific objective and solution. Then, we compute the average across all objectives.
Let $n$ be the number of solutions and $m$ the number of objectives. Let $X$ be an $n\times m$ matrix where $x_{i,j}$ is the train performance for the $j^{th}$ objective of the $i^{th}$ solution, with $1\leq i\leq n$ and $1\leq j\leq m$. Similarly, let $X^{\prime}$ be an $n\times m$ matrix where $x^{\prime}_{i,j}$ is the test performance for the $j^{th}$ objective of the $i^{th}$ solution. Let $H_{\iota}$ be the HV computed on the train performance $X$. $\partial H_{\iota}/{\partial x_{i,j}}$ is the partial derivative of the HV with respect to $x_{i,j}$. We define the Pareto delta function ($P_{\Delta}$) as
2.2 DOSA-MO: Dual-stage Optimizer for Systematic overestimation Adjustment in Multi-Objective problems
The DOSA-MO is a wrapper algorithm that implements a two-phased approach. Initially, regression models are trained to predict overestimation based on selected solutions´ characteristics: their the original fitness, its variance and the feature set size in our case study. The original fitness and its variance are general characteristics for solutions to any ML problem, while the feature set size is specific for feature selection solutions. Then, in the subsequent phase, these regression models are used to refine the MO optimizer’s solution assessment, focusing on better model selection to enhance overall performance. In more details, DOSA-MO consists of three main steps.
@@ 2.2 DOSA-MO: Dual-stage Optimizer for Systematic overestimation Adjustment in Multi-Objective problems
1. Generating solution sets for overestimation prediction. It collects solutions to be used as training samples to learn how to adjust the objective functions that are used to evaluate solutions for supervised learning tasks, such as classification accuracy. This consists in running MO optimizers in a k-fold CV loop. For each fold, a solution set is produced using only training data, and its performance is measured on the left-out samples.
@@
2. Training of regression models for overestimation. For each objective DOSA-MO trains a regression model on the samples collected during step 1. Each sample contains as independent variables three meta-features of the solutions that are potentially predictive of the overestimation. They are the original fitness (i.e. the fitness used by the optimizer to choose the best solutions, measured using only training data, applying inner k-fold CV in our case study), the standard deviation of that fitness (we measure it using bootstrap), and the number of features included in the solution (number of genes forming the biomarker in our case study). The dependent variable is the overestimation: the difference between the original fitness and the fitness that is computed on new data through CV. Not all solutions can be considered as equally important. We might expect solutions that are in crowded areas of the Pareto front to be selected less often by a decision maker. To take this into account each sample is weighted according to the partial derivative of the HV (constructed using the original fitnesses) with respect to the considered solution and objective. The weights for each fold and objective are scaled to sum to 1. A regression model is trained on these samples for each objective. We minimize the absolute error, when allowed by the specific model, since the impact of an error on the HV is approximately linear for small errors. 3. Generating the solution set using adjusted performance. A second MO optimizer is deployed to generate a solution set (each solution refers to a feature set in our case study), with objectives that are systematically adjusted by previously trained regression models for overestimation. In the revised process, each original objective function is replaced by a pipeline that initially computes the function’s result, its standard deviation, and the feature count of the solution. This data is then feed to the corresponding adjuster regression model, which predicts the overestimation. The final adjusted performance is calculated by deducting the predicted overestimation from the original fitness value. The solution set generated by the MO optimizer during this final run represents the output of the whole DOSA-MO.
2.2.1 Pseudocode formulation of the DOSA-MO algorithm

### 29 Feb 2024 v3

# Dual-stage optimizer for systematic overestimation adjustment applied to multi-objective genetic algorithms for biomarker selection

@@ Abstract
The challenge in biomarker discovery using machine learning from omics data lies in the abundance of molecular features but scarcity of samples. Most feature selection methods in machine learning require evaluating various sets of features (models) to determine the most effective combination. This process, typically conducted using a validation dataset, involves testing different feature sets to optimize the model’s performance. Evaluations have performance estimation error and when the selection involves many models the best ones are almost certainly overestimated. Biomarker identification with feature selection methods can be addressed as a multi-objective problem with trade-offs between predictive ability and parsimony in the number of features. Genetic algorithms are a popular tool for multi-objective optimization but they evolve numerous solutions thus are prone to overestimation. Methods have been proposed to reduce the overestimation after a model has already been selected in single-objective problems, but no algorithm existed capable of reducing the overestimation during the optimization, improving model selection, or applied in the more general multi-objective domain. We propose DOSA-MO, a novel multi-objective optimization wrapper algorithm that learns how the original estimation, its variance, and the feature set size of the solutions predict the overestimation. DOSA-MO adjusts the expectation of the performance during the optimization, improving the composition of the solution set. We verify that DOSA-MO improves the performance of a state-of-the-art genetic algorithm on left-out or external sample sets, when predicting cancer subtypes and/or patient overall survival, using three transcriptomics datasets for kidney and breast cancer.
Data and source code availability. The original gene expression data used in this study is from public repositories. The preprocessed data, source code, and detailed numerical results are available in a public server (github.com/UEFBiomedicalInformaticsLab/BIODAI/tree/main/DOSA_MO).
@@ 1 Introduction
Molecular biomarker discovery with machine learning (ML) is usually limited by data that includes many features but few samples [1]. This renders trained models prone to overfitting and the evaluation prone to estimation error. Hyperparameter tuning, including feature selection, is often a crucial aspect of the optimization process where the model’s performance is assessed using a training, validation, and test paradigm. This essential aspect involves choosing the most effective subset of features, essentially optimizing the ML’s model. Selecting the best model from many often leads to overestimation, given by a significant gap between validation set performance and actual, real-world performance, a phenomenon known as the “winner’s curse”. The models that fit the noise present in the validation set are advantaged, a phenomena sometimes referred as overfitting on the validation set [2]. This is also exacerbated by data scarcity. Seeking higher accuracies by expanding hyperparameter configurations can enhance model performance on validation sets. However, this often results in heightened overestimation, leading to reduced or potentially negative impact on test set performance.
@@
In biomarker discovery, the focus is often on optimizing the accuracy of machine learning models using the selected molecular features, while also minimizing the number of features to ensure clinical feasibility and resource efficiency. Characterising all the best compromises (or trade-offs) between predictive value and feature set size is a multi-objective (MO) optimization problem [3, 4], and it can be solved by means of MO feature selection (MOFS) techniques [5, 6, 7, 8]. These techniques aim to identify not just a single best solution, as in single-objective (SO) problems, but rather a Pareto front of solutions. This front is the set of optimal solutions that illustrate the trade-offs between different objectives. However, all candidate solutions (or biomarker models selected by the employed MOFS technique) are evaluated on the validation set, which can result in the overestimation of the performance of the selected models.
K-fold cross-validation (CV) stands as the predominant methodology for machine learning assessment, with its advantages and limitations extensively explored, particularly in SO scenarios [9, 10, 11]. K-fold CV returns a model trained on all the available samples and an estimation of its performance computed by averaging $k$ CV results. The obtained model performance is usually underestimated when only one hyperparameter configuration is used. However, it tends to be overestimated when multiple configurations are evaluated, and the model (or feature set) with the best performance is returned [2]. This situation becomes particularly pronounced in MO approaches, where identifying multiple favorable trade-offs often leads to an increased number of models to be evaluated. A practical, illustrative example is presented in Cattelani et al. Fig. 1, where it is evaluated the performance of various gene expression-based molecular feature sets for classifying breast cancer subtypes [7]. In this MO setting, balanced accuracy and feature set size serve as evaluation metrics. The expected balanced accuracy tends to increase with the addition of more features. Similarly, overestimation – quantified as the difference between training and test set performance (the gap between “inner cv” and “test” results in the figures) – also grows as the estimation improves. These correlations may suggest that characteristics of the solutions might be predictive of the degree of overestimation.
Although various methods have been developed to enhance performance estimation in model selection using k-fold CV, their design and implementation have been limited to SO problems. Tsamardinos et al. [12] compared double CV, the Tibshirani and Tibshirani method [13], and nested CV in their ability to improve the estimation of the fitness for SO problems. These algorithms modify fitness estimations but do not change the model selection process; the chosen model remains the same as it would be using simple hyperparameter optimization with k-fold CV for model evaluation. Automated ML (AutoML) tools offer an approach designed to explore various model and hyperparameter combinations. These tools aim to identify and deliver the most effective model along with an assessment of its performance. In Tsamardinos et al. six AutoML tools were compared [14]. Of these, only one had a predictive performance estimation strategy that could adjusts for multiple model validations (limitedly to SO problems and not affecting model selection), while most of the tools have the necessity to withold a test set for an unbiased estimation of the performance of the winning model, thus loosing samples from the final model training.
No significant efforts have been directed toward improving performance estimation in MO problems, of which SO problems are a special case and which are more relevant and applicable to biomarker discovery in high-dimensional omics data. Moreover, while there are approaches for SO problems to enhance the performance estimation of a chosen model, these models are still selected based on unadjusted estimations. Consequently, this does not lead to any improvement in the actual model selection process. On the other hand, strategies that involve subtracting a constant value from model evaluation metrics, used for ranking multiple solutions, fall also short. This approach alone is insufficient to significantly impact the ranking order and thus remains ineffective for enhancing the model selection process. However, if the adjustment takes into account some characteristic of the models, such as the variance in performance during the validation phase or the feature set size, it might change their order. It may be argued that such an adjustment could improve the model selection if the model ranking would use the adjusted estimations. In summary, to the best of our knowledge, no previous work experimented the effectiveness of methods for mitigating the overestimation in MO problems using ML algorithms. Additionally, no previous work applied the adjustment to the performance estimation during the optimization of the solution set, thus affecting the selection.
Figure 1: Sequence of operations for external validation of DOSA-MO. A MO problem is defined with multiple objectives, e.g. cancer subtype classification, prediction of survival, and parsimony in the feature set size. A dataset (e.g. TCGA breast omics data) is fed to DOSA-MO for its optimization process in 3 steps. In step 1 it performs a k-fold CV with a wrapped MO optimizer (e.g. the NSGA3-CHS GA) and collects the solutions from all the folds. From each solution and objective a sample is constructed. It has the fitness expected by the wrapped MO, its SD, and the feature set size as independent variables, the overestimation (expected fitness minus fitness assessed on left-out set) as dependent variable, and the partial derivative on the HV with respect to this fitness measurement as sample weight. In step 2 these samples are used to train regression models for overestimation, and new adjusted objective functions are created. In step 3 a wrapped MO optimizer is run with the adjusted objective functions, impacting how the models are selected. The solution set is the output of DOSA-MO (e.g. a set of biomarkers). It might be beneficial to use a faster wrapped optimizer in step 1 than in step 3 since step 1 uses k-fold CV. An external dataset (e.g. SCAN-B) is used to externally validate the solution set.
Here we present the DOSA-MO (Dual-stage Optimizer for Systematic overestimation Adjustment in Multi-Objective problems), a cutting-edge algorithm aimed at predicting and adjusting for overestimation in MO problems. Initially, a wrapped MO optimizer is integrated with ML algorithms, to create a series of preliminary solutions (feature sets in our case study). These initial solutions are then used to train regression models that are specifically designed to predict performance overestimation. This prediction is based on characteristics of the solutions, such as the variance in evaluation metrics (for instance, the variance of balanced accuracy in internal validations) or the size of the feature set itself. As wrapped MO optimizer for our case study we used a modification of the Non-dominated Sorting Genetic Algorithm III (NSGA3) [15]: NSGA3 with Clone-Handling method and Specialized mutation operator (NSGA3-CHS, see Section 2.3). Subsequently, the algorithm executes a wrapped MO optimizer once more, but with a key distinction: it now utilizes the regression models to deliver adjusted fitness evaluations, enhancing the model selection process. The final resulting models are then trained on all the available samples, thus no samples are lost in order to compute their expected performance on new data. In our benchmarking study, which concentrates on selecting gene expression-based feature sets for cancer subtype classification and patient survival prediction, we have empirically demonstrated that DOSA-MO effectively mitigates overestimation and improves model selection. It consistently delivers improved performance estimations in biomarker discovery across diverse population-based cohort datasets. Additionally, two novel measures are introduced in our study: MO Performance Error (MOPE) and Pareto Delta ($P_{\Delta}$). To the best of our knowledge, these are the first to be designed to evaluate the discrepancy between the performance expected by the algorithm and the actual performance observed on new samples in MO problems. These assessments consider the entirety of the solution set determined by the MO optimizer, providing a comprehensive view of the algorithm’s effectiveness. The flow of a DOSA-MO use case is depicted in Fig. 1.
Disambiguations for the most technical terms can be found in Supplementary Section 3.
@@ 2 Methods
2.1 DOSA-MO: algorithm description
The DOSA-MO algorithm wraps a MO optimizer, serving two purposes: improve the estimate of the solutions’ performance, and increase the overall performance of the solution set (set of biomarker models in our case study). In our case study, these optimizers are represented by genetic algorithms (GAs) [16] specifically designed for MO problems and paired with supervised ML algorithms, such as the Gaussian Naïve Bayes (NB) classifier to distinguish cancer subtypes or the Cox Proportional-Hazards Model (Cox) for survival analysis. In more details, DOSA-MO consists of three steps.
@@ 2.1 DOSA-MO: algorithm description
1. Generating solution sets for overestimation prediction. It collects solutions to be used as training samples to learn how to adjust the objective functions that are used to evaluate solutions, such as the classification accuracy. This consists of running MO optimizers in a k-fold CV loop. For each fold, a solution set is produced using only training data, and its performance is measured on the left-out samples.
@@
2. Training of regression models for overestimation. For each objective DOSA-MO trains a regression model on the samples collected during step 1. Each sample contains as independent variables three meta-features of the solutions that are potentially predictive of the overestimation. They are the original fitness (i.e. the fitness used by the optimizer to choose the best solutions, measured using only training data, applying inner k-fold CV in our case study), the standard deviation (SD) of that fitness (we measure it using bootstrap), and the number of features included in the solution (number of genes forming the biomarker in our case study). The dependent variable is the overestimation: the difference between the original fitness and the fitness computed on new data through CV. Solutions cannot be considered as equally important. We might expect solutions that are in crowded areas of the non-dominated front to be selected less often by a decision maker. Consequently, each sample is weighted according to the partial derivative of the HyperVolume (HV, constructed using the original fitnesses) [17] with respect to the considered solution and objective. The weights for each fold and objective are scaled to sum to 1. We minimize the absolute error, when allowed by the specific regression model, since the impact of an error on the HV is approximately linear for small errors. 3. Generating the solution set using adjusted performance. A second MO optimizer is deployed to generate a solution set (each solution refers to a feature set in our case study), with objectives that are adjusted by previously trained regression models for overestimation. Each original objective function is replaced by a pipeline that initially computes the function’s result, its SD, and the feature count of the solution. This data is then feed to the corresponding adjuster regression model, which predicts the overestimation. The final adjusted performance is calculated by deducting the predicted overestimation from the original fitness value. The solution set generated by the MO optimizer during this final run represents the output of the whole DOSA-MO.
Depending on the MO optimizers used, step 1 might be computationally expensive. A strategy to limit its cost when using GAs is described in Supplementary Section 1.3.
Step 2 includes computing the SD of the original fitness. Our method for doing that combining bootstrap with k-fold CV is described in Supplementary Section 1.4.
2.2 Pseudocode formulation of the DOSA-MO algorithm